OpenStudy (anonymous):

Solve for R 2r+5=3.5r+2

OpenStudy (mathlegend):

Okay, well this is just like any other equation. So lets get the 'r' variable to one side of the equation. If I were you I would keep the 'r' variable positive.

OpenStudy (mathlegend):

So what will you subtract first.. 2r or 3.5r?

OpenStudy (mathlegend):

@tutu show me what you will subtract.

OpenStudy (anonymous):

Umm,the 3.5??

OpenStudy (mathlegend):

Why that? It will make a negative number. Lets go with the 2.

OpenStudy (mathlegend):

2r+5=3.5r+2 -2r -2r ------------ 5=3.5r+2

OpenStudy (mathlegend):

So now show me the next step.

OpenStudy (anonymous):

Ummm -2 on both sides?

OpenStudy (mathlegend):

Good so we have... 3=1.5r

OpenStudy (mathlegend):

r = ?

OpenStudy (anonymous):

Yeah then divide and you get 0.5

OpenStudy (mathlegend):

No. 1.5 + 1.5 = 3 or 1.5 * 2 = 3

OpenStudy (mathlegend):

So what does r = ?

OpenStudy (mathlegend):

@tutu do you understand?

OpenStudy (anonymous):

\[2r+5=3.5r+2\] \[5=1.5r+2\] \[3=1.5r\] \[2=r\]

OpenStudy (anonymous):

Now I do